OpenStudy (anonymous):

Simplify the expression: sin 2a/cos a

OpenStudy (anonymous):

(2(sin a) (cos a))/(cos a) = 2 * sin a

OpenStudy (anonymous):

*Double angle formula if you look in your book or notes, is what was just used in case you want to use it again:)

OpenStudy (anonymous):

KK. Thanks!

OpenStudy (anonymous):

Use algebra to prove the identity: sin t/1-cos t = 1+cos t/ sin t